OpenStudy (anonymous):

find exact value of cot(-17pi/4)

OpenStudy (raden):

cot(-17pi/4) = - cot(17pi/4) = - cot(16pi/4 + pi/4) = - cot(4pi + pi/4) = - cot(2*2pi + pi/4) = - cot(pi/4) = - 1/tan(pi/4) i sure, u have remembered the value of tan(pi/4), just plug it to there